A telephone booth 210 CM tall casts a shadow 600cm long. At the same time, a nearby fire hydrant casts a shadow 240cm long. Find the height of the fire hydrant.
1
Expert's answer
2020-11-01T17:41:09-0500

Let the height of the fire hydrant = "x"

height of telephone booth "\div" telephone booth shadow length = "x\div" fire hydrant shadow length

"\frac{210}{600} = \frac{x}{240}"

"600x = 50400"

"x=\frac{50400}{600}=84cm" (height of the fire quadrant).